OpenStudy (javk):

Can somebody help me understand a transitive relation. I understand both reflexive and symmetric relations, but I am having issues understanding transitive relations.

OpenStudy (michele_laino):

for example let's consider this relation: \[a|b\], namely a divides b, or a is a divisor of b

OpenStudy (michele_laino):

which means that exists a integer k such that b = a*k

OpenStudy (michele_laino):

I can state that \[a|b\] is an equivalence relation

OpenStudy (javk):

ok

OpenStudy (michele_laino):

now we can check the transitive property as follows: if \[a|b\quad {\text{and}}\quad b|c\]

OpenStudy (michele_laino):

then I state that \[a|c\]

OpenStudy (michele_laino):

here is the proof: by definition, we know that exists an integer k such that: b=k*a furthermore, by definition, I know that exists an integer h, such that: c=h*b

OpenStudy (michele_laino):

so we can write this: \[c = hb = h\left( {ka} \right) = \left( {hk} \right)a\]

OpenStudy (michele_laino):

since hk is again an integer number, then we can can say that c divides a, namely the thesis

OpenStudy (michele_laino):

that is an example of transitive property

OpenStudy (javk):

so does that mean that there is a relationship that connects every three members?

OpenStudy (michele_laino):

better is to say an implication, namely: \[{\text{if }}\left( {a|b\quad {\text{and}}\quad b|c} \right)\quad {\text{then: }}a|c\]

OpenStudy (javk):

but say for the sake of understanding...I'm trying to visualize it...is it true?

OpenStudy (michele_laino):

yes!

OpenStudy (javk):

ok, got it thanks

OpenStudy (michele_laino):

thanks!